# Data-driven Radiative Magnetohydrodynamics Simulations with the MURaM Code: the Emergence of Active Region 11158 and the X2.2 Flare

### v1

@@ 3.2 Magnetic Structure and Plasma Dynamics of the Eruption
Figure 2: Course of the eruption. Panel (a) shows the horizontally averaged density from the combined data of the evo. and flare runs, which illustrate the slow rise and eruption of a plasma-hosting magnetic flux rope. The overlay black line plots the synthesized GOES 1-8Å soft X-ray flux, which indicates a flare above X class. Panels (b)–(g) present 3D visualizations of the plasma and magnetic field structures of the eruption. The corresponding time stamps are marked by the white dashed lines in Panel (a). The opacity of the features is chosen on the basis of the the plasma density in the erupted flux rope, whereas lower and higher values are made transparent. The magnetic field lines are calculated from seed points that are randomly distributed in the coronal volume above the sunspots. The color of the density features and fieldlines reflects the plasma temperature on a logarithmic scale.
@@
3.2.1 Slow Rise and Fast Eruption of the Flux Rope
@@
We use density structures to trace the kinematic evolution of the flux rope, as in our previous studies (Wang et al., 2023) and in observations where ejected plasma is tracked (Cheng et al., 2020, and reference therein). Figure 2(a) presents the horizontally averaged plasma density as a function of time. The 3000 s time period covers the slow rise of the flux rope before the rapid eruption. The height $z$ starting from 4 Mm is displayed on a logarithmic scale to highlight the lower part of the domain.
@@
We calculate the soft X-ray flux with a temperature response function of the GOES-15 satellite, which is overlaid on the density profile. The peak flux of $1.97\times 10^{-4}$ W m-2 is found at $t\approx 300$ s, which corresponds to an X2 class flare. The similarity in the soft Xray flux is very likely only a coincidence, because real flares undergo physical processes, e.g., those related to nonthermal particles (Aschwanden et al., 2016; Warmuth & Mann, 2016), that are beyond the capability of the current numerical model. Therefore, the actual value of the soft X-ray flux is only compared with other simulations under similar approximations and is considered as an indicator of the strength of energy release and plasma heating. This flare is thus far the largest event, compared with early MURaM simulations in the literatures that produced flares corresponding to lower M (Chen et al., 2023b; Rempel et al., 2023) or C class (Cheung et al., 2019), with magnetic energy release on the order of $10^{31}$ erg.
@@
For more 1000 s before the eruption, the flux rope steadily rises at a very small speed of approximately 2 km/s and progresses to rapid eruption around $t=250$ s during the impulsive phase of the flare. The lead edge in the density profile is a piston-driven shock (presented later with more details). The main body of the flux rope rises at an average speed on the order of 200 km/s. Although the domain height is the maximum that our computational resources can afford, the shock and ejected plasma reaches the top boundary at approximately $t=420$ s and $t=700$ s, respectively. The top boundary is open for outflows but may still affect the upward propagation of the CME. Observations (e.g., Figure 9 in Schrijver et al., 2011) found a radial expansion front between 40–100 Mm above the surface with a deduce speed of 300 km/s at 4 min after the onset, which is similar to the propagation in this simulation.
3.2.2 Large-scale Magnetic Structure in the Eruption
@@
We briefly present the magnetic and plasma structures in the early stage of the eruption in Panels (b) – (g) of Figure 2. Before eruption, the flux rope can be found along the PIL between N1 and P2 and is accompanied by sheared arcades on both sides. Cool plasma is hosted in the dip formed by a twisted magnetic field. Panel (c) shows the snapshot in the impulsive phase of the flare, during which reconnections occur between the pre-existing flux rope with ambient arcades and gives rise to 10 MK plasma in the core area between N1 and P2, which is the location of observed X-ray sources Zharkov et al. (2011); Wang et al. (2012). Panel (d) displays a snapshot shortly after the flare peak. The two hot arcades connecting to from the inner spots N1 and P2 to the outer spots P1 and N2, respectively, now reconnect to form a long sigmoid, which is considered as a new large flux rope that roots in P1 and N2, and at the same moment, post-flare loops connecting N2 and P1 are formed below the reconnection site.
@@
The snapshots in the gradual phase are presented in Panels (e) – (g) of Figure 2. The long sigmoid continues to rise and push the plasma upward with ongoing reconnection below, which is similar to the mechanism found in more generalized numerical experiments (Jiang et al., 2021). New post-flare loops continue to form left (solar east or lower $x$) to right (solar west or higher $x$). The propagation of reconnection is supposed to drive the extension of flare ribbons, which is discussed later in this paper. Panel (g) illustrates the reconnetion between the envelope field on the two sides of the erupted flux rope and the formation of longer and higher post-flare loops that are filled with plasma over 10 MK.
3.2.3 Plasma Dynamics in the Large-scale Corona
@@
We analyze the emission measured weighted velocity $v_{\rm LOS}$, which represents the Doppler velocity of a spectral line forming in the given temperature range, as done in our previous studies (Chen et al., 2023a; Chen, 2025). Figure 3 presents $v_{\rm LOS}$ in three temperature ranges that capture the chromosphere and transition region plasma, warm coronal plasma of million K, and hottest plasma around and above 10 MK.
@@
The second row displays Doppler maps in the gradual phase of the flare. The downflow ribbons significantly extend along the PIL, as we expected from the development of magnetic reconnection as shown in Figure 3(e) and (f). Moreover, the eruption produces more spatially extensive upflows of a few hundred km/s. Fan et al. (2024) reported a significant fieldline expansion at similar locations, which may explain the dimming observed during this eruption (Dissauer et al., 2018).
The bottom row presents a snapshot when the atmosphere is recovering from the highly dynamic state caused by the eruption. The time scale might not be realistic because the real eruption impacts a much larger volume than the domain size of the simulation and lasts significantly longer. Nevertheless, we see that, during this stage, the plasma velocities at all temperatures decrease to their typical values during a quiescent period.
Last but not least, the upward eruption of the magnetic structure during the impulsive and peak phases of the flare is accompanied by strong horizontal expansion as well, which eventually leads to a downward push at the outer rim of the active region, as shown by the increasing and sharpening downflow rims in Panels (b) and (e) of Figure 3. This gives rise to an intriguing interactions between different layers of the atmosphere as presented in detail in the following section.
3.3 CME Driven Shock and the Chromospheric Moreton Wave
@@
We present in Figure 4 (a) the vertical velocity $v_{z}$ near the peak of the flare, which is the identical snapshot as Figure 2 (d) and the top row of Figure 3 in a yzy-z plane that cuts through the core region of the eruption. The plasma after the piston-driven shock has an upward velocity of 500 km/s, in contrast to the mild 200 km/s downward velocity in the background. In comparison, Afanasyev et al. (2023) and Fan et al. (2024) reported velocities above or close to 1000 km/s after the shock front. The highest speed in this vertical slice is found immediately above the reconnection site at $z\approx 20$ Mm, which is up to 1500 km/s and accompanied by a 500 km/s downflow right below, which appears as a red point in Figure 4 (a).
@@
Panels (b) and (c) display two time-distance diagrams of $v_{z}$ along a vertical and a horizontal slit that are placed at locations indicated by the two dashed lines in Figure 4 (a), respectively. Panel (b) shows that the upward propagating disturbance is set off in the impulsive phase at approximately $t=270$ s. The velocity gradient increases and develops into a shock (when it is resolved by 5 grid points in the vertical direction) at $t\approx 320$ s and $z\approx 170$ Mm. Afterward, the shock propagates at a rather constant speed of approximately 1000 km/s. Moreover, Figure 4 (b) shows similar branches, as illustrated in Figure 2 (a), where we see a fast propagating shock front and a flux rope that rise at a small speed.
@@
The diagram in Figure 4 (c) aims to capture the impact of the eruption at the outer part of the active region. We find numerous thin ridges that extend from the spine, which is the upward erupting flux rope, to both sides. The slopes of the ridges indicate a propagation speeds on the order of 104 km/s. After $t=275$ s, a few ridges (on the lower $y$ side) exhibit downflows, which means that the atmosphere above may exert a downward push to that below this layer and those even deeper.
@@
Similar behaviors were also reported by Afanasyev et al. (2023) and Fan et al. (2024). Unfortunately, their model may not have a setup that allows for studying the dynamics in the lower atmosphere. Therefore, we examine this effect in our simulation in a deep horizontal layer at 2 Mm height in the chromosphere, as shown in Figure 4 (d). In the outer rim of the active region, which is dominated by downflows, a belt of slightly increased downward velocities (e.g, at $y=60$ Mm) can be discerned. This disturbance is better revealed in Panel (e) by a time-distance diagram along the $y$-slit placed at the exact location of the yzy-z plane. After $t=275$ s, we see in this deep chromosphere layer a propagating downflow wave front to the south, which will generate a signal on the red-wing of chromospheric spectral lines such as how Moreton waves were discovered (Moreton & Ramsey, 1960). The propagation speed of approximately 1600 km/s is surprisingly large compared with the typical sound speed in the chromosphere. It follows the long-known argument that such chromospheric waves cannot be sustained in the chromosphere but rather imprint the impact of waves/shocks propagating in the corona (Uchida, 1968).
@@
Panel (f) of Figure 4 displays a time-distance diagram along the $x$-slit marked by the horizontal dashed line in Panel (d). The imprint of coronal waves/shocks is only found to the east of the active region, and together with Figure 4 (e), it clearly demonstrates the asymmetric propagation of the Moreton wave.
3.4 Formation of Flare Ribbons
3.4.1 Plasma Dynamics in Flare Ribbons
@@
Figure 5: The flare ribbons at $t=410$ s. Only the lowermost 12.8 Mm (200 grid points) are considered in this analysis to highlight the emission in the lowest part of the domain and to exclude the contribution from ejected chromospheric plasma with the erupted flux rope. The first row displays emission measures binned to three wide temperature ranges. The second row shows the vertical component of field-aligned velocities, i.e., the upward evaporation flows, at the surface of a certain temperature, as indicated above Panels (d)–(f). Panels (g)–(i) present the vertical acceleration driven by the pressure gradient force along the magnetic field. Panels (j)–(l) illustrate the thermal conduction flux at these three temperature surfaces.
@@
Instead of the line-of-sight integrated velocity shown in Figure 3, we illustrate the plasma dynamics in flare ribbons via the velocity in at a constant temperature surface. Panels (d)–(f) reveal chromospheric evaporations and condensations in the flaring atmosphere in three layers through the field-aligned vertical velocity $v_{\rm b_{z}}$, which is given by
@@
in the same temperature surfaces in Panels (g) – (i).
@@
Strong downflows are found in temperatures over one MK at locations where the brightest ribbons are observed in the center of the active region and are cospatial with the strongest downward pressure gradient acceleration. Slightly weaker condensations are present in the extended ribbons. In the hottest surface, which corresponds to a slightly higher geometric height, the partition of upflows prevails, whereas are some points along the ribbon may still exhibit downflows. An extensive area off the flare ribbons is dominated by upflows from one MK. This is also consistent with the blue patches in Figure 5 (h), where the pressure gradient drives a very weak condensation at lower temperatures but evident evaporation at higher temperatures.
@@
The dynamics observed in the simulation are attributed to the energy deposition in the chromosphere. In this simulation, conduction is the main agent that transports the magnetic energy that is released by reconnections in the corona down to the lower atmosphere. The bottom row of Figure 5 displays the thermal conduction flux in the three layers with increasing temperature (increasing height as well). A downward energy flux greater than $10^{11}$ erg/cm2/s is presents in the higher layers, which is comparable to the strongest energy injection rate assumed in radiative hydrodynamics calculations for solar flares (e.g., Allred et al., 2015; Reep et al., 2015; Kowalski et al., 2017), whereas the energy flux decays significantly with height and becomes the order of $10^{9}$ erg/cm2/s at the surface of lower temperatures. This decay indicates a deposition of energy in this volume, particularly along the flare ribbons, where significant condensations and evaporations are driven by downward/upward pressure gradient forces.
@@
We mark the locations of the sun-quake sources found by Kosovichev (2011) and Zharkov et al. (2011) during this flare in Panels (d) and (k). The observed sun-quake sources, which should be in a further deeper layer, are more or less cospatial with points of strong condensations, i.e., a high energy injection rate. Moreover, two of them appear to be the footpoints of the pre-existing flux rope along the central PIL, and the middle one is right beneath the reconnection site in the early stage of the flare, as illustrated in Figure 2. Although the present model does not account for all the physics needed to model the trigger of the phenomenon, the comparison suggests a nontrivial similarity between the energy deposition site in the simulation and that in the real flare.
3.4.2 Evolution of the flare ribbons and Coronal Magnetic Reconnection

### v2

@@ III.2 Magnetic Structure and Plasma Dynamics of the Eruption
Figure 2: Course of the eruption. Panel (a) shows the horizontally averaged density from the combined data of the evo. and flare runs, which illustrate the slow rise and eruption of a plasma-hosting magnetic flux rope. The overlay black line plots the synthesized GOES 1-8Å soft X-ray flux, which indicates a flare above X class. Panels (b)–(g) present 3D visualizations of the plasma and magnetic field structures of the eruption. The corresponding time stamps are marked by the white dashed lines in Panel (a). The opacity of the features is chosen on the basis of the plasma density in the erupted flux rope, whereas lower and higher values are made transparent. The magnetic field lines are calculated from seed points that are randomly distributed in the coronal volume above the sunspots. The color of the density features and fieldlines reflects the plasma temperature on a logarithmic scale.
@@
III.2.1 Slow Rise and Fast Eruption of the Flux Rope
@@
We use density structures to trace the kinematic evolution of the flux rope, as in our previous studies (C. Wang et al., 2023) and in observations where ejected plasma is tracked (X. Cheng et al., 2020, and reference therein). Figure 2(a) presents the horizontally averaged plasma density as a function of time. The 3000 s time period covers the slow rise of the flux rope before the rapid eruption. The height $z$ starting from 4 Mm is displayed on a logarithmic scale to highlight the lower part of the domain.
@@
We calculate the soft X-ray flux with a temperature response function of the GOES-15 satellite, which is overlaid on the density profile. The peak flux of $1.97\times 10^{-4}$ W m-2 is found at $t\approx 300$ s, which corresponds to an X2 class flare and is close to the X2.2 class of the real flare (C. J. Schrijver et al., 2011). The similarity in the soft X-ray flux is very likely only a coincidence because real flares undergo physical processes, e.g., those related to nonthermal particles (M. J. Aschwanden et al., 2016; A. Warmuth & G. Mann, 2016), that are beyond the capability of the current numerical model. Therefore, the actual value of the soft X-ray flux is only compared with other simulations under similar approximations and is considered an indicator of the strength of energy release and plasma heating. This flare is thus far the largest event, compared with early MURaM simulations in the literature that produced flares corresponding to lower M (F. Chen et al., 2023b; M. Rempel et al., 2023) or C class (M. C. M. Cheung et al., 2019), with magnetic energy release on the order of $10^{31}$ erg.
@@
For more than 1000 s before the eruption, the flux rope steadily rises at a very small speed of approximately 2 km/s and progresses to rapid eruption around $t=250$ s during the impulsive phase of the flare. The lead edge in the density profile is a piston-driven shock (presented later with more details in Section III.3). The main body of the flux rope rises at an average speed on the order of 200 km/s. Although the domain height is the maximum that our computational resources can afford, the shock and ejected plasma reaches the top boundary at approximately $t=420$ s and $t=700$ s, respectively. The top boundary is open for outflows but may still affect the upward propagation of the CME. Observations (e.g., Figure 9 in C. J. Schrijver et al., 2011) revealed a radial expansion front between 40–100 Mm above the solar surface with a deduce speed of 300 km/s at 4 min after the onset, which is similar to the propagation in this simulation.
III.2.2 Large-scale Magnetic Structure in the Eruption
@@
We briefly present the magnetic and plasma structures in the early stage of the eruption in Panels (b) – (g) of Figure 2. Before eruption, the flux rope can be found along the PIL between N1 and P2 and is accompanied by sheared arcades on both sides. Cool plasma is hosted in the dip formed by a twisted magnetic field. Panel (c) shows the snapshot in the impulsive phase of the flare, during which reconnections occur between the pre-existing flux rope with ambient arcades and gives rise to 10 MK plasma in the core area between N1 and P2, which is the location of the observed X-ray sources S. Zharkov et al. (2011); S. Wang et al. (2012). Panel (d) displays a snapshot shortly after the flare peak. The two hot arcades connecting from the inner spots N1 and P2 to the outer spots P1 and N2, respectively, now reconnect to form a long sigmoid, which is considered as a new large flux rope that roots in P1 and N2, and at the same moment, post-flare loops connecting N2 and P1 are formed below the reconnection site.
@@
The snapshots in the gradual phase are presented in Panels (e) – (g) of Figure 2. The long sigmoid continues to rise and push the plasma upward with ongoing reconnection below, which is similar to the mechanism found in more generalized numerical experiments (C. Jiang et al., 2021). New post-flare loops continue to form left (solar east or lower $x$) to right (solar west or higher $x$). The propagation of reconnection is supposed to drive the extension of flare ribbons, which is discussed later in this paper. Panel (g) illustrates the reconnection between the envelope field on the two sides of the erupted flux rope and the formation of longer and higher post-flare loops that are filled with plasma over 10 MK.
III.2.3 Plasma Dynamics in the Large-scale Corona
@@
We analyze the emission measured weighted velocity $v_{\rm LOS}$, which represents the Doppler velocity of a spectral line forming in the given temperature range, as in our previous studies (F. Chen et al., 2023a; F. Chen, 2025). $v_{\rm LOS}$ within a given temperature range $T_{\rm L},T_{\rm H}]$ is calculated as
$v_{\rm LOS}=\frac{\sum\limits_{T_{i}=T_{\rm L}}^{T_{\rm H}}v_{z}(\log T_{i}){\rm EM}_{z}(\log T_{i})}{\sum\limits_{T_{i}=T_{\rm L}}^{T_{\rm H}}{\rm EM}_{z}(\log T_{i})}.$ (6)
${\rm EM}_{z}(\log T_{i})$ is a standard MURaM output for analysis, as described in F. Chen et al. (2022), which records the emission measures integrated along a given axis (the vertical axis $z$ here) for temperatures from $\log T$=4.5 to the highest value in the domain with an interval of $\Delta\log T=0.1$. Similarly, $v_{z}(\log T_{i})$ is the line-of-sight average velocity for the same temperature discretization of ${\rm EM}_{z}$. Moreover, this average uses the emission measure as a weight, when the contribution of each cell in a vertical column is counted into the corresponding temperature interval. Figure 3 presents $v_{\rm LOS}$ in three temperature ranges that capture the chromosphere and transition region plasma, warm coronal plasma of million K, and hottest plasma around and above 10 MK.
@@
The second row displays Doppler maps in the gradual phase of the flare. The downflow ribbons significantly extend along the PIL, as we expected from the development of magnetic reconnection, as shown in Figure 3(e) and (f). Moreover, the eruption produces more spatially extensive upflows of a few hundred km/s. Y. Fan et al. (2024) reported a significant fieldline expansion at similar locations, which may explain the dimming observed during this eruption (K. Dissauer et al., 2018).
The bottom row presents a snapshot when the atmosphere is recovering from the highly dynamic state caused by the eruption. The time scale might not be realistic because the real eruption impacts a much larger volume than the domain size of the simulation and lasts significantly longer. Nevertheless, during this stage, the plasma velocities at all temperatures decrease to their typical values during a quiescent period.
Last but not least, the upward eruption of the magnetic structure during the impulsive and peak phases of the flare is accompanied by strong horizontal expansion as well, which eventually leads to a downward push at the outer rim of the active region, as shown by the increasing and sharpening downflow rims in Panels (b) and (e) of Figure 3. This gives rise to intriguing interactions between different layers of the atmosphere, as presented in detail in the following section.
III.3 CME Driven Shock and the Chromospheric Moreton Wave
@@
In Figure 4 (a), we present the vertical velocity $v_{z}$ in a yzy-z plane that cuts through the core region of the eruption near the peak of the flare, which is the same time as Figure 2 (d) and the top row of Figure 3. The plasma after the piston-driven shock has an upward velocity of 500 km/s, in contrast to the mild 200 km/s downward velocity in the background. In comparison, A. N. Afanasyev et al. (2023) and Y. Fan et al. (2024) reported velocities above or close to 1000 km/s after the shock front. The highest speed in this vertical slice is found immediately above the reconnection site at $z\approx 20$ Mm, which is up to 1500 km/s and accompanied by a 500 km/s downflow right below, which appears as a red point in Figure 4 (a).
@@
Panels (b) and (c) display two time-distance diagrams of $v_{z}$ along a vertical and a horizontal slit that are placed at locations indicated by the two dashed lines in Figure 4 (a), respectively. Panel (b) shows that the upward-propagating disturbance is set off in the impulsive phase at approximately $t=270$ s. The velocity gradient increases and develops to a shock (when it is resolved by 5 grid points in the vertical direction) at $t\approx 320$ s and $z\approx 170$ Mm. Afterward, the shock propagates at a rather constant speed of approximately 1000 km/s. Moreover, Figure 4 (b) shows similar branches, as illustrated in Figure 2 (a), where we see a fast propagating shock front and a flux rope that rise at a small speed.
@@
The diagram in Figure 4 (c) aims to capture the impact of the eruption at the outer part of the active region. We find numerous thin ridges that extend from the spine, which is the upward erupting flux rope, to both sides. The slopes of the ridges indicate a propagation speed on the order of 104 km/s. After $t=275$ s, a few ridges (on the lower $y$ side) exhibit downflows, which means that the atmosphere above may exert a downward push to that below this layer and even deeper.
@@
Similar behaviors were also reported by A. N. Afanasyev et al. (2023) and Y. Fan et al. (2024). Unfortunately, their model may not have a setup that allows for studying the dynamics in the lower atmosphere. Therefore, we examine this effect in our simulation in a deep horizontal layer at 2 Mm height in the chromosphere, as shown in Figure 4 (d). In the outer rim of the active region, which is dominated by downflows, a belt of slightly increased downward velocities (e.g, at $y=60$ Mm) can be discerned. This disturbance is better revealed in Panel (e) by a time-distance diagram along the $y$-slit placed at the exact location of the yzy-z plane. After $t=275$ s, we see in this deep chromosphere layer a propagating downflow wave front to the south, which will generate a signal on the red wing of chromospheric spectral lines such as how Moreton waves were discovered (G. E. Moreton & H. E. Ramsey, 1960). The propagation speed of approximately 1600 km/s is surprisingly large compared with the typical sound speed in the chromosphere. It follows the long-known argument that such chromospheric waves cannot be sustained in the chromosphere but rather imprint the impact of waves/shocks propagating in the corona (Y. Uchida, 1968).
@@
Panel (f) of Figure 4 displays a time-distance diagram along the $x$-slit marked by the horizontal dashed line in Panel (d). The imprint of coronal waves/shocks is only found east of the active region, and together with Figure 4 (e), it clearly demonstrates the asymmetric propagation of the Moreton wave.
III.4 Formation of Flare Ribbons
III.4.1 Plasma Dynamics in Flare Ribbons
@@
Figure 5: The flare ribbons at $t=410$ s. Only the lowermost 12.8 Mm (200 grid points) are considered in this analysis to highlight the emission in the lowest part of the domain and to exclude the contribution from ejected chromospheric plasma with the erupted flux rope. The first row displays emission measures binned to three wide temperature ranges. The second row shows the vertical component of field-aligned velocities, i.e., the upward evaporation flows, on the surface at a certain temperature, as indicated above Panels (d)–(f). Panels (g)–(i) present the vertical acceleration driven by the pressure gradient force along the magnetic field. Panels (j)–(l) illustrate the thermal conduction flux on these three temperature surfaces.
@@
Instead of the line-of-sight integrated velocity shown in Figure 3, we illustrate the plasma dynamics in flare ribbons via the velocity on a constant temperature surface. Such surfaces are naturally 3D structures in space. We use a linear interpolation on $\log T$ to obtain the height $z$ of a constant temperature surface as an xyx-y map and interpolate column-wise from 3D cubes of quantities of interest to extract the values on this temperature isosurface.
Panels (d)–(f) reveal chromospheric evaporations and condensations in the flaring atmosphere in three layers through the field-aligned vertical velocity $v_{\rm b_{z}}$, which is given by
@@
on the same temperature surfaces in Panels (g) – (i).
@@
Strong downflows are found in temperatures over one MK at locations where the brightest ribbons are observed in the center of the active region and are cospatial with the strongest downward pressure gradient acceleration. Slightly weaker condensations are present in the extended ribbons. On the hottest surface, which corresponds to a slightly higher geometric height, the partition of upflows prevails, whereas some points along the ribbon may still exhibit downflows. An extensive area off the flare ribbons is dominated by upflows from one MK. This is also consistent with the blue patches in Figure 5 (h), where the pressure gradient drives very weak condensations at lower temperatures but evident evaporations at higher temperatures.
@@
The dynamics observed in the simulation are attributed to energy deposition in the chromosphere. In this simulation, conduction is the main agent that transports the magnetic energy that is released by reconnections in the corona down to the lower atmosphere. The bottom row of Figure 5 displays the thermal conduction flux in the three layers with increasing temperature (increasing height as well). A downward energy flux greater than $10^{11}$ erg/cm2/s is present in the higher layers, which is comparable to the strongest energy injection rate assumed in radiative hydrodynamics calculations for solar flares (e.g., J. C. Allred et al., 2015; J. W. Reep et al., 2015; A. F. Kowalski et al., 2017), whereas the energy flux decays significantly with height and becomes the order of $10^{9}$ erg/cm2/s on the surface at lower temperatures. This decay indicates a deposition of energy in this volume, particularly along the flare ribbons, where significant condensations and evaporations are driven by downward/upward pressure gradient forces.
@@
We mark the locations of the sun-quake sources found by A. G. Kosovichev (2011) and S. Zharkov et al. (2011) during this flare in Panels (d) and (k). The observed sun-quake sources, which should be in a further deeper layer, are more or less cospatial with points of strong condensations, i.e., a high energy injection rate. Moreover, two of them appear to be the footpoints of the pre-existing flux rope along the central PIL, and the middle one is right beneath the reconnection site in the early stage of the flare, as illustrated in Figure 2. Although the present model does not account for all the physics needed to model the trigger of the phenomenon, the comparison suggests a nontrivial similarity between the energy deposition site in the simulation and that in the real flare.
III.4.2 Evolution of the flare ribbons and Coronal Magnetic Reconnection